# CMB parity asymmetry from unitary quantum gravitational physics

### v2

@@ IV CMB parity asymmetry
the Fourier coefficients $a_{\ell m}$ are characterized by the angular power spectrum $C_{\ell}$. Because $\mathcal{T}$ (and $\mathcal{P}\mathcal{T})$ is broken by the slow-roll parameters in Eq.18, we expect the quantum field during inflation to evolve time asymmetrically at parity conjugate points in physical space. This leads to an angular power spectrum $C_{\ell}$ of the CMB exhibiting an excess of power in the odd multipoles compared to the even ones, as described by Eq.27 in Appendix C.
@@
In Fig. 3, we can visually see the parity asymmetry by the close resemblance between odd-parity temperature maps and the original CMB maps from Planck data (Akrami et al. (2020a)). This reflects that the CMB temperature maps contain an additional pure odd-parity component. The DSI framework exactly generates this additional odd symmetry, as we can see in Eq.27. It is worth noticing here that if we take the average of even and odd power spectra, we recover the well-known near-scale invariant power spectrum of Standard Inflation (SI). The benefit of DSI is that there are no additional free parameters. The parity asymmetry only depends on the spectral index $n_{s}\approx 0.9634\pm 0.0048$ on the pivot scale $k_{\ast}=0.05\,{\rm Mpc}^{-1}$ ($\ell\approx 800$) as measured by Planck data (Akrami et al. (2020b)).
@@
In Fig. 5 we present how the observational data significantly prefer DSI over SI. Observed quantities include the observed low quadrupolar amplitude $C_{2}=C_{\ell=2}$, the even-to-odd multipole ratio of angular power spectra:
@@
and the mean $Z^{1}=<Z(\hat{n})>$ of the parity map $Z(\hat{n})\equiv\mathcal{T}(\hat{n})\mathcal{T}(-\hat{n})$ displayed in Fig. 4, where $-\hat{n}=(\pi-\theta,\pi+\varphi)$ represents the parity conjugate or antipode of the sky position $\hat{n}=(\theta,\varphi)$. The likelihood probability associated with DSI in these plots is up to 650 times larger than that of SI. DSI favors odd parity, leading to a lower $C_{2}$. This, in turn, diminishes the significance of other CMB anomalies (Schwarz et al. (2016)), such as the Hemispherical Power Asymmetry, the axis of Evil, and the anomalous 2-point correlation function (see Gaztañaga and Kumar (2024) for further details).
Figure 4: Parity map: $Z(\hat{n})=\mathcal{T}(\hat{n})\mathcal{T}(-\hat{n})$ for Fig. 3 with galactic mask. This is a $\mathcal{P}$ conjugate image of itself. Positive values (red) correspond to the symmetric (even parity) component of $\mathcal{T}$, while negative values (blue) give the antisymmetric (odd parity) component. There is a clear excess of negative values, indicating a preference for odd parity. This, together with Fig. 3 serves as a visual testament to the unitary QFTCS in nature.
Figure 5: Parity measurements derived from different mask (component separation) CMB temperature maps $\mathcal{T}(\hat{n})$ (dashed lines) against data simulations (grey areas represent 1, 2, and 3 sigma contours), alongside predictions from standard inflationary (SI) models (in red) and Direct-Sum inflation (DSI) (in blue). The later is 650 more probable.
@@ V Conclusions
In this paper, we underscore the pivotal role of unitary Quantum Field Theory in Curved Spacetime (QFTCS) in elucidating the intricacies of early Universe physics. Contrary to prevailing notions, we contend that the issue of unitarity in QFTCS is not inherently tied to quantum gravity at Planck scales, but rather stems from throwing away the $\mathcal{P}\mathcal{T}$ symmetry for quantum fields in curved spacetime.
With the $\mathcal{P}\mathcal{T}$-symmetric formulation of Schrödinger equation Eq.1, the DQFT framework reinstates unitarity, which was previously thought to be lost in curved spacetime. Fig. 1-2 visually represent this concept, illustrating how a hypothetical classical observer consistently experiences pure states without encountering information loss beyond the horizon. Furthermore, the striking correlation illustrated in Fig. 3-5 between the observed hot and cold mirrored structures of the CMB map at antipodal points strongly supports the unitary treatment of inflationary quantum fluctuations within the framework of DQFT. Our findings indicate that superhorizon fluctuations at antipodes manifest as $\mathcal{P}\mathcal{T}$ conjugates, a prediction derived from direct-sum QFTCS. This prediction is substantially favored with a probability over 650 times greater than that of the standard QTFCS.
@@ Acknowledgements.
EG acknowledges grants from Spain Plan Nacional (PGC2018-102021-B-100) and Maria de Maeztu (CEX2020-001058-M). KSK acknowledges the Royal Society for the Newton International Fellowship.
Appendix A Klein-Gordon field operator in DQFT
The Klein-Gordon field operator for DQFT in Minkowski spacetime split into direct-sum
@@ Appendix A Klein-Gordon field operator in DQFT
which defines the positive norm states in the direct-sum vacuums $|0_{T}\rangle=|0_{+}\rangle\oplus|0_{-}\rangle=\begin{pmatrix}|0_{+}\rangle\\ |0_{-}\rangle\end{pmatrix},\,a_{(\pm)}|0_{\pm}\rangle=0$. The second condition in Eq.23 implies commutativity of operators $\left[\hat{\phi}_{+},\,\hat{\phi}_{-}\right]=0$, which can be seen as a new causality condition. The The Fock space in DQFT is the direct-sum of two Fock spaces (geometric superselection sectors) $\mathcal{F}_{T}=\mathcal{F}_{+}\oplus\mathcal{F}_{-}$ corresponding to the components of field operators ϕ^±subscript^italic-$\hat{\phi}_{\pm}$ that lead to the quantum states forward and backward in time at parity conjugate regions of position space. These field operators are defined through the pairs of creation and annihilation operators Eq.23. Since the Minkowski spacetime is $\mathcal{P}\mathcal{T}$ symmetric, the scalar field operator in DQFT respects the same symmetry. Generalizing this structure to all quantum fields in Minkowski spacetime, such as complex scalar, vector, and fermionic degrees of freedom, is very straightforward. Any standard model (single particle) state, whether it is particle ($|\rm SM\rangle)$ or anti-particle ($|\overline{\rm SM}\rangle$, is expressed in DQFT as direct-sum of two components
@@
with respect to $\mathcal{P}\mathcal{T}$ conjugate splitting of the vacuum $|0_{M}\rangle=|0_{M+}\rangle\oplus|0_{M-}\rangle$. For further details on DQFT in particular, the CPT invariance of scattering amplitudes which holds in both geometric superselection sectors, and quantizing other fields of the standard model, see Kumar and Marto (2025).
Appendix B Appendix B: DQFT in dS
Applying DQFT in dS (Gaztañaga and Kumar (2024); Kumar and Marto (2025)), the canonically rescaled scalar field ($\phi\to e^{Ht}\phi$) operator is
$\hat{\phi}$ $=\frac{1}{\sqrt{2}}\hat{\phi}_{+}\left(\tau,\,\textbf{x}\right)\oplus\frac{1}{\sqrt{2}}\hat{\phi}_{-}\left(-\tau\,-\textbf{x}\right)\,,$ (25)

### v4

@@ IV.1 CMB temperature sky fluctuations
where $k_{s}=7\times 10^{-5}\,{\rm Mpc}^{-1}$ and $j_{\ell}\left(z\right)$ are spherical Bessel functions of the first kind. Because the time-reversal symmetry $\mathcal{T}$ is broken by the slow-roll parameters in Eq. 28, we expect the quantum field during inflation to evolve asymmetrically in time at parity-conjugate points in physical space (resulting in Eq.27 and Eq. 30). Following Eq. 30 and Eq. 36, we obtain an angular power spectrum $C_{\ell}$ of the CMB that exhibits an excess of power in odd multipoles compared to even ones, as described below in Eq.38.
In Fig. 3, we can visually identify this parity asymmetry by the strong resemblance between the odd-parity component of the temperature map and the original CMB maps from Planck data Akrami et al. (2020b). This reflects that the CMB temperature maps contain an additional pure odd-parity component, as indicated by Eq. 36. The DSI framework naturally generates this additional odd symmetry.
@@
It is worth noting that if we take the average of even and odd power spectra, we recover the well-known near scale-invariant power spectrum of standard inflation (SI) with $\Delta\mathcal{P}_{v}\simeq 0$. The advantage of the DSI model is that it requires no additional free parameters: the parity asymmetry depends solely on the spectral index $n_{s}\approx 0.9634\pm 0.0048$ at the pivot scale $k_{\ast}=0.05\,{\rm Mpc}^{-1}$ ($\ell\approx 800$), as measured by Planck Akrami et al. (2020a). Inflationary quantum fluctuations in DSI are non-markovian (See Sec. 5.4 of Gaztañaga and Kumar (2024)). So compute here the scale-dependent features of power spectra for the large wavelength modes that affect the $-\ell$ features of the CMB angular power spectrum.
IV.2 DSI vs SI Model Comparison and Simulations
To assess the significance of the observed parity asymmetry in the CMB, we compare $10^{6}$ simulated realizations of the data under two models: the SI model and the DSI model. Specifically, we evaluate the posterior probability $p(M|D)$ of each model $M$ given the data $D$.
This approach contrasts with the standard practice in the CMB community, which typically estimates the likelihood $p(D|M)$—the probability of the data given a specific model—to assess the significance of low multipole anomalies. However, this likelihood-based method presupposes the model and can lead to inflated uncertainties, especially because the $\Lambda$CDM model with SI generally predicts more large-scale power than is observed. This mismatch increases the sampling variance, thereby reducing the apparent significance of observed parity anomalies. For example, the low measured quadrupole $C_{2}$ has $p(D|M)=2.62\%$ while the posterior value is 29 times smaller: $p(M|D)=0.09\%$, as shown in Table 1.
Both the data and simulations are processed in the same way. Full details of the simulation setup are provided in Appendix D of Gaztañaga and Kumar (2024). Here we focus on the comparison of the $10^{6}$ all-sky-masked realizations of the measured $C_{\ell}$ power spectrum. We used the $C_{\ell}$ values measured by the Planck 2018 analysis. We consider four Planck component-separated map estimates: Commander (COMMm), SMICA (SMICm), SEVEM (SEVEm), and NILC (NILCm), each analyzed at $N_{\rm side}=2048$. The corresponding sky fractions are $f_{\rm sky}=88.8\%$, $84.2\%$, $83.8\%$, and $78.6\%$, respectively. The appropriate mask is applied to each simulation set. We use the HEALPix software package [^5] https://healpix.sourceforge.io/ for all-sky map analysis and visualizations. The SI model corresponds to the Planck 2018 best fit scale invariant $\Lambda$ CDM angular power spectrum, denoted $C_{\ell}^{\rm SI}$ (shown as a black dashed line in the top panel of Fig. 4) and goes throught the middle of the $C_{\ell}$ measurements (cyan line and shaded region 68%), as expected. The DSI model introduces a parity-violating modulation to the SI spectrum as follows:
$C_{\ell}^{\rm DSI}=C_{\ell}^{\rm SI}\left[1+(-1)^{\ell+1}\Delta\mathcal{C}_{\ell}\right]$ (38)
where the fractional modulation $\Delta\mathcal{C}_{\ell}$ is:
$\Delta\mathcal{C}_{\ell}=\frac{1}{C_{\ell}^{\rm SI}}\int_{0}^{k_{c}}\frac{dk}{k}A_{s}\left(\frac{k}{k_{s}}\right)^{n_{s}-1}j_{\ell}^{2}\left(\frac{k}{k_{s}}\right)\Delta\mathcal{P}_{v}(k)$ (39)
Here, $\Delta\mathcal{P}_{v}(k)$ is given in Eq. 33, and the standard $\Lambda$CDM angular power spectrum is (see Eq.37) :
$C_{\ell}^{\rm SI}=\int_{0}^{\infty}\frac{dk}{k}A_{s}\left(\frac{k}{k_{\ast}}\right)^{n_{s}-1}j_{\ell}^{2}\left(\frac{k}{k_{s}}\right)$ (40)
The SI prediction corresponds to $\Delta\mathcal{P}_{v}=0$, while the DSI one is given by Eq.33. A cut-scale $k_{c}=0.02k_{\ast}$ is used in Eq.39 as Eq.33 is accurate enough for $-\ell$ or large angular scales. This cutoff scale is also related to the coarse-graining scale of stochastic inflation (Gaztañaga and Kumar (2024)), which determines the large wavelength modes that have already become classical on the onset of inflation while the mode $k_{\ast}$ exits the horizon. In DSI, these large-wavelength modes create parity-asymmetric inhomogeneities that correct the FLRW spacetime on large scales. The parity asymmetry ceases to exist towards the small scales because the $\mathcal{P}\mathcal{T}$ symmetry is recovered towards the short-distance scales. This can be seen in Fig. 11 of Gaztañaga and Kumar (2024). An intuitive explanation for this is that large wavelength modes experience quantum gravitational effects more than those of short wavelength modes. For the small scales $k>k_{c}$, the predictions of DSI become almost the same as with the (near) scale-invariant power spectrum as contributions from $\Delta\mathcal{P}_{v}$ term become negligible.
The resulting modulation in Eq. 38 oscillates with amplitude $\sim 20\%$ for multipoles $\ell<10$, producing a measurable parity asymmetry (see dashed red line in top panel of Fig. 4). We generate $10^{6}$ Monte Carlo realizations for both SI and DSI models, allowing a direct and statistically robust comparison with the Planck data.
@@ IV.2 DSI vs SI Model Comparison and Simulations
In Fig. 4 and Table 1 we present how the posterior probabilities from the $10^{6}$ realizations of the observational data very significantly prefer the mean DSI model prediction over the SI model. We compare the quadrupole amplitude $C_{2}=C_{\ell=2}$, the even-to-odd multipole ratio of angular power spectra:
@@
[…]
Parity SI SI DSI ratio
indicator  $p[D|M]$  $p[M|D]$  $p[M|D]$   DSI/SI
$C_{2}$ 2.62 % 0.09 % 3.3 % 37
$R^{TT}$ 1.0 % 0.7 % 39.5 % 56
$Z^{1}$ 3.89 % 1.12 % 45.3 % 40
$C_{2},R^{TT}$ 0.12 % 0.003 % 1.96 % 653
$Z^{1},R^{TT}$ 0.45 % 34.6 % 77
$Z^{1},C_{2}$ 0.016 % 2.65 % 166
[…]
Figure 5: TOP: Parity map $Z(\hat{n})=\mathcal{T}(\hat{n})\,\mathcal{T}(-\hat{n})$ derived from the SMICA map shown in Fig. 3, including the Planck galactic mask. This map is invariant under the parity operator $\mathcal{P}$, and encodes the degree of symmetry between temperature values at antipodal sky positions. Positive regions (red) indicate even-parity (symmetric) structure in the CMB temperature field $\mathcal{T}$, while negative regions (blue) highlight odd-parity (antisymmetric) structure. The dominance of negative values visually indicates a preference for odd parity in the observed sky. Notably, the average value $\langle Z\rangle\simeq-0.1$ supports the DSI prediction, whereas the SI model predicts $\langle Z\rangle\simeq 0.1$ (see also Fig. 4). BOTTOM: Relative deviation (in percent) between the observed distribution of $Z(\hat{n})$ values (for both unmasked SMICA, black, and masked SMICA, red) and the mean SI prediction (black dashed line). The data show a clear excess of negative $Z$ values and a deficit of positive ones, compared to expectations under SI. The shaded cyan band represents the 68% confidence interval from $10^{6}$ SI simulations. The red line shows the DSI model (labeled DQFT), which offers a significantly better match to the observed distribution. Together with Fig. 3, this provides visual and statistical evidence for the underlying parity-violating quantum field theory in curved spacetime (QFTCS) responsible for the observed asymmetry.
@@ V Conclusions
In this paper, we underscore the pivotal role of unitary Quantum Field Theory in Curved Spacetime (QFTCS) in elucidating the intricacies of early Universe physics. Contrary to prevailing notions, we contend that the issue of unitarity in QFTCS is not inherently tied to quantum gravity at Planck scales, but rather stems from throwing away the vital role of $\mathcal{P}\mathcal{T}$ transformations in the construction of vacuum for quantum fields in curved spacetime. In the context of de Sitter spacetime, which is $\mathcal{P}\mathcal{T}$ symmetric manifold, we illustrated that one can restore unitarity of QFTCS by taking into consideration geometrical aspects of spacetime.
With the $\mathcal{P}\mathcal{T}$ based formulation of the Schrödinger equation Eq. 1, we embrace a new understanding of quantum theory where a single quantum state is expressed as direct-sum of two components that describe the parity conjugate regions of physical space with opposite arrows of time. The DQFT framework constructs quantum fields with opposite arrows of time at parity conjugate regions of physical space. This new geometric construction of quantum fields echoes with discrete (a)symmetries of the spacetime manifold. Within this new description, the spacetime horizon acts as a $\mathcal{P}\mathcal{T}$ mirror which contains any information within the horizon in the form of pure states. This reinstates unitarity in de Sitter spacetime, which was previously thought to be lost in standard frameworks of quantization. Fig. 12 visually represent this concept, illustrating how a hypothetical classical observer consistently experiences pure states without encountering information loss beyond the horizon. The DQFT framework can be applied to any curved spacetime without any necessity of $\mathcal{P}\mathcal{T}$ symmetry in the manifold. The construction of DQFT only prescribes that every quantum field should be expressed via two components describing parity conjugate regions of physical space with opposite arrows of time, where the quantum vacuum is a geometrical direct-sum based on $\mathcal{P}\mathcal{T}$ operations. With this construction, quantum fields could evolve asymmetrically at parity-conjugate points if the time reversal symmetry is broken. Thus, in manifolds like de Sitter spacetime, where the manifold is $\mathcal{P}\mathcal{T}$ symmetric, quantum fields geometrically obey the symmetry and lead to equal two-point correlations at parity conjugate regions of physical space.
Applying the framework of unitary QFTCS to inflationary quantum fluctuations, we derived a prediction for a parity asymmetry in the CMB originating from time-reversal symmetry breaking during inflation. It is well-known that time is a parameter in quantum theory. If one considers quantum fluctuations in gravitational physics, all the time-dependent quantities that describe the dynamics of spacetime act as parameters. Thus, in quantum gravity (even at the linearized level, which is the case of inflationary quantum fluctuations), time reversibility is much more complex than a naive thinking of time coordinate reflection. In our construction, the slow-roll parameters that describe inflationary spacetime dynamics break the $\mathcal{P}\mathcal{T}$ symmetry of the de Sitter vacuum in DQFT, leading to parity asymmetric evolution of the inflationary quantum fluctuations. We have shown that this leads to non-trivial correlation between anti-podal points ($\left(\theta,\,\varphi\right)$ and $\left(\pi-\theta,\,\pi+\varphi\right)$) of the CMB sky. The striking correlation illustrated in Fig. 35 between the observed hot and cold mirrored structures of the CMB map at antipodal points provides strong support for the unitary treatment of inflationary fluctuations within the DQFT framework.
Our findings indicate that superhorizon fluctuations at antipodes manifest as $\mathcal{P}\mathcal{T}$ conjugates, a prediction derived from direct-sum QFTCS (i.e., the direct-sum inflation, DSI). This prediction is strongly supported by the data, with a posterior probability over 650 times greater than that of standard inflation, which is based on non-unitary QFTCS (see Table 1). The observed parity asymmetry also provides a unified explanation for other large-scale CMB anomalies, such as the so-called hemispherical power asymmetry Gaztañaga and Kumar (2024).
To assess the significance of the low-multipole anomalies, we evaluate the posterior probability $p(M|D)$ of each model $M$ given the data $D$. This Bayesian approach contrasts with the standard practice in the CMB community, which evaluates $p(D|M)$—the likelihood of the data under the SI model. For example, the low measured quadrupole $C_{2}$ has $p(D|M)=2.62\%$ while the posterior value is 29 times smaller: $p(M|D)=0.09\%$, as shown in Table 1. The former approach tends to inflate uncertainties due to the excess power predicted by $\Lambda$CDM at large scales, which increases sampling variance and thereby reduces the apparent significance of observed parity asymmetry.
Our study highlights the importance of QFTCS and its non-trivial role in underpinning the quantum aspects of gravity. By sticking to the foundational aspects of QFTCS, we addressed the long-standing issues associated with the understanding of CMB anomalies. Our grounded approach not only shed new light on understanding quantum gravity but also allowed us to derive predictions that necessarily addressed the observations. This approach complements and poses new challenges to the past investigations Ashtekar et al. (2021); Kitazawa and Sagnotti (2015) that addressed CMB anomalies as evidence of yet unknown frameworks of Planck scale quantum gravity.
@@ Acknowledgements.
EG acknowledges grants from the Spanish Plan Nacional (PGC2018-102021-B-100) and Maria de Maeztu (CEX2020-001058-M). KSK acknowledges the Royal Society for the Newton International Fellowship.
Appendix A Quantum harmonic oscillator in direct-sum quantum mechanics
The purpose of this section is to explicitly illustrate the application of direct-sum Schrödinger to the context of the simple harmonic oscillator whose Hamiltonian is given by
^$\hat{\mathbb{H}}=\frac{E}{2}\left(\hat{p}^{2}+\hat{x}^{2}\right)=\underbrace{\frac{E}{2}\left(\hat{p}_{+}^{2}+\hat{x}_{+}^{2}\right)}_{\hat{\mathbb{H}}_{+}}\oplus\underbrace{\frac{E}{2}\left(\hat{p}_{-}^{2}+\hat{x}_{-}^{2}\right)}_{\hat{\mathbb{H}}_{-}}=\begin{pmatrix}\hat{\mathbb{H}}_{+}&&0\\ 0&&\hat{\mathbb{H}}_{-}\end{pmatrix}$ (42)
Following Eq. II, we express the operators $\hat{x}_{\pm}$ and $\hat{p}_{\pm}$ in terms of the two pairs of creation and annihilation operators $\left(a,\,a^{\dagger}\right)$ and $\left(b,\,b^{\dagger}\right)$ associated with the sectorial Hilbert spaces $\mathcal{H}_{\pm}$ as
$\hat{x}_{+}$ $=\sqrt{\frac{1}{2}}\left(a+a^{\dagger}\right),\quad\hat{p}_{+}=-i\frac{d}{dx_{+}}=\frac{i}{\sqrt{2}}\left(a^{\dagger}-a\right)$ (43)
$\hat{x}_{-}$ $=\sqrt{\frac{1}{2}}\left(b+b^{\dagger}\right),\quad\hat{p}_{-}=i\frac{d}{dx_{-}}=-\frac{i}{\sqrt{2}}\left(b^{\dagger}-b\right)$
with
$\Big{[}\hat{x}_{+},\,\hat{p}_{+}\Big{]}$ $=i,\quad\Big{[}\hat{x}_{-},\,\hat{p}_{-}\Big{]}=-i$ (44)
$\Big{[}a,\,a^{\dagger}\Big{]}$ $=\Big{[}b,\,b^{\dagger}\Big{]}=1,\quad\Big{[}a,\,b^{\dagger}\Big{]}=\Big{[}a,\,b\Big{]}=0\,.$
Solving the direct-sum Schrödinger equation Eq. 1 for the direct-sum split harmonic oscillator Hamiltonian operator Eq. 42 yields the wave function $\Psi_{n}\left(x\right)$ for the different energies $E_{n}=\frac{1}{2}\omega\left(n+\frac{1}{2}\right)$ as
$\Psi_{n}(x)=\langle x|\Psi_{n}\rangle\equiv\begin{cases}\frac{1}{\sqrt{2^{n+1}n!}}\left(\frac{1}{\pi}\right)^{1/4}e^{-\frac{1}{2}x_{+}^{2}}H_{n}\left(x_{+}\right)e^{-iE_{n}t_{p}},\quad x_{+}\gtrsim 0\\ \frac{1}{\sqrt{2^{n+1}n!}}\left(\frac{1}{\pi}\right)^{1/4}e^{-\frac{1}{2}x_{-}^{2}}H_{n}\left(x_{-}\right)e^{iE_{n}t_{p}},\quad x_{-}\lesssim 0\end{cases}$ (45)
where we followed Eq. 7. It is important to recall the description of quantum mechanics with two components $|\Psi\rangle$ of the parity conjugate regions with geometric superselection sectors $\mathcal{H}_{\pm}$, which reinstate the $\mathcal{P}\mathcal{T}$ symmetry of the physical system respected at the quantum mechanical level. In Fig. 6 we present the schematic description of the quantum harmonic oscillator in direct-sum QM.
Figure 6: This represents the schematic understanding of the quantum harmonic oscillator in direct-sum quantum mechanics in which a single quantum state is written as direct-sum of two components $|\Psi\rangle=\frac{1}{\sqrt{2}}\left(|\Psi_{+}\rangle\oplus|\Psi_{-}\rangle\right)$ described by geometric superselection sector Hilbert spaces $\mathcal{H}_{\pm}$ (the total Hilbert space that represents the state $|\Psi\rangle$ that describe the parity conjugate regions of physical space where arrow of times are oppositely defined. Both $|\Psi_{\pm}\rangle$ are the positive energy states in the respective sectorial Hilbert spaces.
Appendix B Klein-Gordon field operator in DQFT
The Klein-Gordon field operator for DQFT in Minkowski spacetime $ds^{2}=-dt_{m}^{2}+d\textbf{x}^{2}$ split into direct-sum
@@ Appendix B Klein-Gordon field operator in DQFT
which defines the positive norm states in the direct-sum Minkowski vacuums
$|0_{M}\rangle=|0_{M_{+}}\rangle\oplus|0_{M_{-}}\rangle=\begin{pmatrix}|0_{M+}\rangle\\ |0_{M_{-}}\rangle\end{pmatrix},\quad a_{(\pm)}|0_{\pm}\rangle=0.$ (49)
The second condition in Eq.48 implies commutativity of operators $\left[\hat{\phi}_{+},\,\hat{\phi}_{-}\right]=0$, which can be seen as a new causality condition. The The Fock space in DQFT is the direct-sum of two Fock spaces (geometric superselection sectors) $\mathcal{F}_{T}=\mathcal{F}_{+}\oplus\mathcal{F}_{-}$ corresponding to the components of field operators ϕ^±subscript^italic-$\hat{\phi}_{\pm}$ that lead to the quantum states forward and backward in time at parity conjugate regions of position space. These field operators are defined through the pairs of creation and annihilation operators Eq.48. Since the Minkowski spacetime is $\mathcal{P}\mathcal{T}$ symmetric, the scalar field operator in DQFT respects the same symmetry. Generalizing this structure to all quantum fields in Minkowski spacetime, such as complex scalar, vector, and fermionic degrees of freedom, is very straightforward. Any standard model (single particle) state, whether it is particle ($|\rm SM\rangle)$ or anti-particle ($|\overline{\rm SM}\rangle$, is expressed in DQFT as direct-sum of two components
@@
with respect to $\mathcal{P}\mathcal{T}$ conjugate splitting of the vacuum $|0_{M}\rangle=|0_{M_{+}}\rangle\oplus|0_{M_{-}}\rangle$, where the subscript $M_{\pm}$ mean $\mathcal{P}\mathcal{T}$ conjugate sheets of Minkowski spacetime. Also, we assume the same structure of the geometric superselection rule for all Fock spaces of The SM degrees of freedom i.e., $\mathcal{P}\mathcal{T}$ conjugate regions of spacetime defined the same way for all quantum fields of SM. Note that in each vacuum $|0_{M\pm}\rangle$ a (component) of an anti-particle state can be interpreted as a particle (component) state going backward in time. It means, $|\overline{{\rm SM}}_{+}\rangle$ can be interpreted as $|{\rm SM}_{+}\rangle$ state going backward in time with respect to vacuum $|0_{M+}\rangle$. In the similar way, $|\overline{{\rm SM}}_{-}\rangle$ can be interpreted as $|{\rm SM}_{-}\rangle$ state going backward in time with respect to vacuum $|0_{M-}\rangle$. Thus, Feynman’s interpretation remains the same. Remember that in DQFT, a single quantum state is expressed as direct-sum of two components related by $\mathcal{P}\mathcal{T}$ transformations. To be explicit, the Fermionic field operator in DQFT is quantized as
$\hat{\psi}$ $=\frac{1}{\sqrt{2}}\left(\hat{\psi}_{+}\oplus\hat{\psi}_{-}\right)$ (51)
$\hat{\psi}_{\pm}$ $=\sum_{{\tilde{s}}}\int\frac{d^{3}k}{\left(2\pi\right)^{3/2}\sqrt{2|k_{0}|}}\Bigg{[}c_{{\tilde{s}}(\pm)\textbf{k}}u_{\tilde{s}}(\textbf{k})e^{\pm ik\cdot x}+d_{{\tilde{s}}(\pm)\textbf{k}}^{\dagger}v_{\tilde{s}}(\textbf{k})e^{\mp ik\cdot x}\Bigg{]}$
[…]
$\mathcal{L}_{c}\sim\mathcal{O}_{SM}^{3}=\begin{pmatrix}\mathcal{O}_{SM_{+}}^{3}&0\\ 0&\mathcal{O}_{SM_{-}}^{3}\end{pmatrix}\quad\mathcal{L}_{q}\sim\mathcal{O}_{SM}^{4}=\begin{pmatrix}\mathcal{O}_{SM_{+}}^{4}&0\\ 0&\mathcal{O}_{SM_{-}}^{4}\end{pmatrix}$ (52)
[…]
${\rm CPT}|\rm SM\rangle={\rm CPT}|\rm SM_{+}\rangle\oplus{\rm CPT}|\rm SM_{-}\rangle$ (53)
Appendix C DQFT in dS
Applying DQFT in dS (Gaztañaga and Kumar (2024); Kumar and Marto (2025)), the canonically rescaled scalar field [^7] Here we mean some scalar field $\phi$, not to be confused with the inflaton field ($\phi\to e^{Ht}\phi$) operator is
$\hat{\phi}&=\frac{1}{\sqrt{2}}\hat{\phi}_{+}\left(\tau,\,\textbf{x}\right)\oplus\frac{1}{\sqrt{2}}\hat{\phi}_{-}\left(-\tau\,-\textbf{x}\right)\,,\end{aligned}=\begin{pmatrix}\hat{\phi}_{+}\left(\tau,\,\textbf{x}\right)&&0\\ 0&&\hat{\phi}_{-}\left(-\tau,\,-\textbf{x}\right)\end{pmatrix}$ (54)